OpenStudy (anonymous):

what is the geometric mean of 3 and 15

OpenStudy (anonymous):

Geometric mean of \(a\) and \(b\) is given by: \[G.M. = \sqrt{ab}\] Here a = 3 and b = 15..

OpenStudy (anonymous):

Can you find now??

OpenStudy (anonymous):

@nonna14 you are allowed to reply..

OpenStudy (anonymous):

i know. sorry. um ... i'm not sure how to do that, because 15*3 = 45. and 45 squared is 6.7 ... that's not one of my options

OpenStudy (anonymous):

Are your answers in radical form ??

OpenStudy (anonymous):

\[\sqrt{45} = \sqrt{3 \times 3 \times 5} = ??\]

OpenStudy (anonymous):

so 3 spuared 5 would be the right answer ??

OpenStudy (anonymous):

Yes.. \[\implies 3 \sqrt{5}\]

OpenStudy (anonymous):

ok. thank you so much !

OpenStudy (anonymous):

Welcome dear..